A sample of blood is placed in a centrifuge of radius 16.0 cm. The mass of a red blood cell is 3.0  10−16 kg, and the magnitude of the force acting on it as it settles out of the plasma is 4.0  10−11 N. At how many revolutions per second should the centrifuge be operated?



__ rev/s


1
Expert's answer
2021-10-06T18:16:01-0400

We use the relation between angular velocity, the force of the circular motion, and the acceleration to find the angular velocity "\omega" expressed in rev/s:


"F_c=ma_c=\cfrac{mv^2}{r}=m\omega^2r
\\ \text{ }
\\ \omega= \cfrac{2\pi}{T}=\sqrt{\cfrac{F_c}{mr}}
\\ \implies \omega=\sqrt{\cfrac{4\times10^{-11}\,N}{(3\times10^{-16}\,kg)(0.16\,m)}}
\\ \therefore \omega=912.87\frac{rad}{s}(\frac{1\,rev}{2\pi \,rad})=145.288\frac{rev}{s}"

After substitution, we were able to find that the centrifuge should be operated at least with "\omega =" 145.3 revolutions per second.

Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.